# Shop-R1: Rewarding LLMs to Simulate Human Behavior in Online Shopping via Reinforcement Learning

### v1

@@ 4.1 Experiment Setups
Datasets and models. Our study is built on a proprietary corpus of 52,137 real-world shopping sessions collected from a leading global e–commerce platform. Each session logs the multi-turn interaction between a human customer and the website interface. We enrich each recorded action with a natural language rationale automatically generated by Claude 3.5 Sonnet (see Appendix B for the prompting details). The provided observation context is formatted as simplified HTML [13], which retains essential structural elements while filtering out irrelevant content such as scripts, styling information, and user-specific data. For SFT dataset, we keep each session intact. The model is asked to produce the assistant response, which contains both the rationales and the structured action prediction. For RL dataset, we convert a session into a sequence of <context, action> pairs. The context is the concatenation of (i) all previously observed contexts and (ii) the actions already taken; the target is the next action only. Because every session begins on the home page, there is always at least one observed <context, action> pair before the first prediction step, eliminating the open-world ambiguity of the very first move. To provide the model with slightly richer supervision on the harder behaviours, the two complex actions (click and type_and_submit) each occur about 10% more frequently than the simple terminate action. This mild skew prevents the learner from over-fitting to the trivial case while still maintaining near-uniform coverage, thereby supporting fair and informative per-class evaluation. All experiments fine-tune the publicly available Qwen-2.5-3B-Instruct model. The default 3B parameter backbone offers a favourable compute–performance trade-off.
@@
[…]
@@ 4.2 Experimental Results
Performance comparison with baselines. Main performance comparison results are shown in Tab. 2. Firstly, zero-shot prompting yields low performance: without any task-specific adaptation Qwen-2.5-3B-Instruct achieves only 0.32% exact-action accuracy, confirming that long-horizon web behavior cannot be recovered from generic instruction tuning alone. Second, RL with sparse binary rewards on their own still fail to give the agent meaningful guidance. When we train the policy from scratch under this signal, it reaches only 1.01% exact-match action accuracy and 6.17% type accuracy. Third, a straightforward round of SFT is more effective, boosting performance to 16.76% exact match accuracy and 22.25% type accuracy. This confirms that dense, teacher-forced trajectories are crucial for injecting the structural knowledge (context $\rightarrow$ rationale $\rightarrow$ action) and illustrating the shape of the long-text fields (button labels or search queries) that the binary signal alone cannot convey. Fourth, appending an additional binary-reward RL phase after SFT delivers only mixed results: exact-match action accuracy actually slips to 16.55%, while type-level F1 rises to 28.07%. The agent thus learns to guess the coarse intent better, but it still struggles to reproduce the long-text values that drive the exact-match metric. In other words, the policy becomes better at guessing the coarse action type, but slightly worse at reproducing the fine-grained, long-text values required for an exact match. Lacking finer-grained credit assignment, the binary objective cannot push the model beyond what SFT already achieves and in some respects even pulls it backwards. Training with this objective is furthermore prone to instability and converges substantially more slowly than optimization with richer, structured rewards. Our proposed Shop-R1 framework closes most of this gap. By combining hierarchical rewards, self-certainty signals, format rewards and difficulty-aware scaling, it delivers 27.72% exact-action accuracy (+65% relative to SFT) and pushes action-type accuracy and F1 to 36.40% and 31.28%, respectively. The simultaneous rise of both the coarse (type-level) and fine-grained (exact-match) metrics indicates that Shop-R1 not only identifies the correct intent more often, but also reproduces the long-text values (button labels, text queries) with higher fidelity.
@@
Models Settings Exact Action Acc. Per Action Type Action Type Acc. Per Action Type click type_and_submit terminate click type_and_submit terminate Qwen-2.5-3B-Instruct Zero-shot prompting 0.58% 0.15% 0.00% 38.7% 1.62% 0.00% SFT 4.93% 3.84% 49.80% 8.55% 15.36% 49.80% SFT + RL (Binary) 8.12% 3.25% 45.51% 17.25% 13.88% 45.51% Shop-R1 (Ours) 7.39% 7.53% 81.84% 10.29% 28.66% 81.84% Qwen-1.5-3B-Instruct Zero-shot prompting 1.01% 0.15% 0.39% 10.00% 0.44% 0.39% SFT 4.49% 7.83% 23.44% 15.07% 32.35% 23.44% Shop-R1 (Ours) 3.62% 8.12% 72.85% 6.52% 34.12% 72.85% Qwen-0.5-3B-Instruct Zero-shot prompting 0.43% 0.15% 24.02% 12.90% 4.43% 24.02% SFT 3.19% 7.68% 21.88% 5.94% 26.59% 21.88% Shop-R1 (Ours) 0.72% 3.99% 97.07% 1.01% 17.87% 97.07%
As shown in Tab. 3, we decompose accuracy by action type. Zero-shot prompting shows the classic “intent–content” split. For example, it can guess that a ‘click’ is needed (38.7% type accuracy) yet almost never names the exact UI target (0.58 % exact). Even if SFT can boost the performance but the gain remains uneven, which suggests that teacher forcing alone does not give the model enough credit assignment signal for predicting high-entropy arguments such as search queries. Appending a sparse binary RL phase after SFT still fails to boost these harder text-generation cases. Shop-R1 reshapes those incentives, higher exact match accuracy is achieved, indicating that the model is no longer satisfied with merely selecting correct type but is learning to identify the correct widget and query text as well. To be summarized, dense and structured feedback is essential: it overcomes the no-reward plateau, and makes reward hacking uneconomical.
@@ 4.3 Ablation Study And Analysis
Model size. Tab. 2 and Tab. 3 reveal a consistent scaling trend. In the zero-shot regime, the 3B backbone already outperforms its 1.5B and 0.5B counterparts by a factor of ×45$\times$ 4$\sim$ 5× 4 ∼ 5 on coarse action–type accuracy, confirming that larger models possess stronger out-of-the-box priors for human behavior simulation at the setting of website shopping. After SFT, all sizes gain, yet the improvement is more pronounced for the two smaller backbones than for the 3B model, suggesting that demonstration learning compensates for limited capacity. Shop-R1 lifts every backbone to its best operating point, but the shape of the gains differs by scale. The 3B variant reaches the highest overall numbers while distributing its improvements evenly across the two complicated action types. By contrast, the 0.5B model achieves a comparable headline exact match accuracy (27.72%) almost entirely by over-predicting the easiest ‘terminate’ action (97.07% exact) and largely ignoring the more semantically demanding classes. The 1.5B backbone sits in between, recovering moderate fidelity on ‘click’ and ‘type_and_submit’ while retaining a strong but not overwhelming bias toward ‘terminate’. In short, scaling primarily augments the model’s ability to handle long-text, high-entropy actions; smaller networks can still match aggregate accuracy by exploiting the high-reward termination branch, but they do so at the cost of behavioral diversity. These findings underscore that, although Shop-R1 markedly mitigates capacity limitations, genuine mastery of simulation-oriented web-shopping action prediction tasks continues to benefit from larger backbones.
Sampling temperature. Fig. 2 shows that Shop-R1 is robust to sampling temperature, yet the three evaluation metrics react in distinct ways that reveal how temperature sampling propagates
@@
through the decision hierarchy. Action-type accuracy remains almost constant ($36\%$) across the entire temperature sweep because this metric aggregates all predictions: small mis-classifications in one direction are largely offset by fixes in another, leaving the overall hit rate unchanged. By contrast, the F1 score declines steadily ($31.28\%\rightarrow 28.36\%$) as temperature rises; class-averaging penalizes any asymmetric increase in confusion. Interestingly, a modest boost from the default $\tau=0.6$ to $\tau=0.7$ improves exact-match accuracy to its peak of $28.63\%$: a trace of stochasticity helps the generated response escape local maxima and occasionally assemble the full long-text argument that greedy decoding would miss. When $\tau>0.8$ the added entropy no longer uncovers new correct completions; instead it corrupts fine-grained fields faster than it fixes them, so exact-match plateaus while F1 continues to erode. This pattern is expected since the SFT stage already anchors the model to dataset-specific behavior, privileging faithful simulation over creativity. Taken together, these trends indicate that temperatures in the 0.6–0.8 band offer the best trade-off, preserving robust intent classification, maximizing strict exact-match, and avoiding the metric degradation that emerges once the sampler becomes overly exploratory.
@@
Model Training Scheme Components Exact Action Action Type SFT Format Reward Rationale Reward Reward Scale Action Reward Acc. Acc. F1 Qwen-2.5-3B-Instruct hierarchical 4.63% 36.56% 21.92% hierarchical 2.87% 3.19% 5.04% hierarchical 26.93% 37.25% 33.74% hierarchical 27.83% 27.20% 11.70% binary 27.41% 27.46% 12.11% hierarchical 27.72% 36.40% 31.28%
Training component. Tab. 4 makes clear that every element of Shop-R1 addresses a different pathology. Removing the SFT warm-start cripples the agent: despite having all RL signals, exact-match drops to 4.63%, underscoring that a supervised prior is indispensable for learning the shape of long-text arguments. Omitting the format reward is even more destructive, where exact accuracy plunges to 2.87% and type-level metrics fall below 6% since unparseable JSON outputs earn zero credit, starving the learner of gradient signal. When the self-certainty (rationale) reward is ablated, coarse intent prediction remains strong but exact-match lags the full system by 0.8%, indicating that explicit feedback on the generated rationales mainly tightens the long-text portion of an action rather than its top-level label. Disabling the difficulty-aware reward scaling or reverting to a binary action reward leads to a different failure mode: the model still attains around 27% exact accuracy, yet type-level F1 degrades to 11–12%. Inspection shows that, without either scaling or hierarchical credit, the agent gravitates toward the easy high-reward ‘terminate’ action and rarely ventures into harder ‘click’ or ‘type_and_submit’ cases, which is a classic reward-hacking pattern. The full configuration combines all signals and delivers the best balance demonstrating that each component is necessary: SFT injects linguistic priors, the format reward safeguards parsability, the self-certainty term refines long-text precision, and hierarchical difficulty-scaled rewards prevent degenerate policies while promoting fine-grained action fidelity.
Whole-session v.s. latest-step context. Tab. 5 isolates the impact of including the simplified HTML of each visited page in the action history. Removing this structural cue slashes exact-match accuracy from 27.72% to 14.74%, a nearly 50% relative loss, while coarse action-type accuracy drops more modestly. The sharp divergence indicates that, although the model can still infer which action type of interaction is likely next from the dialogue trace alone, it struggles to generate the fine-grained arguments, the precise button label or query string, without access to the page’s detailed context. Interestingly, the class-balanced F1 score rises slightly, suggesting that the only latest-step context
Table 5: Comparison of model performance when using either the whole-session context or only the latest-step as input.
Settings Exact Action Action Type Acc. Acc. F1 whole-session 27.72% 36.40% 31.28% latest-step 14.74% 30.46% 33.48%
@@ 5 Conclusion
Appendix A System Prompt

### v2

@@ 4.1 Experiment Setups
Datasets and models. Our study is built on Shop-CART dataset consisting of a corpus of 52,137 real‐world shopping sessions collected from a leading global e–commerce service (Lu et al., 2025b). Each session logs the multi‐turn interaction between a human customer and the website interface. More dataset details can be found in Appx. C. We enrich each recorded action with a natural language rationale automatically generated by Claude 3.5 Sonnet (see Appx. D for the prompting details). The provided observation context is formatted as simplified HTML (Lu et al., 2025e), which retains essential structural elements while filtering out irrelevant content such as scripts, styling information, and user-specific data. For SFT dataset, we keep each session intact. The model is asked to produce the assistant response, which contains both the rationales and the structured action prediction. For RL dataset, we convert a session into a sequence of <context, action> pairs. The context is the concatenation of (i) all previously observed contexts and (ii) the actions already taken; the target is the next action only. Because every session begins on the home page, there is always at least one observed <context, action> pair before the first prediction step, eliminating the open‐world ambiguity of the very first move. To provide the model with slightly richer supervision on the harder behaviors, the two complex actions (click and type_and_submit) each occur about 10% more frequently than the simple terminate action. This mild skew prevents the learner from over-fitting to the trivial case while still maintaining near-uniform coverage, thereby supporting fair and informative per-class evaluation. All experiments fine‐tune the publicly available Qwen-2.5-3B-Instruct model. The default 3B parameter backbone offers a favourable compute–performance trade‐off.
@@
Training setups. Our codebase is built on verl (Sheng et al., 2024), and all experiments were conducted on NVIDIA A100 GPUs (80 GB). We leveraged Fully Sharded Data Parallelism (FSDP) in PyTorch (Zhao et al., 2023) to maximize training efficiency. The default policy optimization algorithm is Group Relative Policy Optimization (GRPO) (Shao et al., 2024). Input sequences were padded or truncated to a maximum context length of $32\text{k}$ tokens, and the default sampling temperature is $0.6$. We set the per-device batch size to 1, yielding a global batch size of 64. For supervised fine-tuning (SFT) we trained for 4 epochs with a learning rate of $2\times 10^{-5}$; for reinforcement learning (RL) we trained for 500 steps with a learning rate of $1\times 10^{-7}$. By default, we set the DARS factor to 1000, and use $\alpha=0.005$ and $\beta=0.001$ to weight the corresponding reward terms.
@@ 4.2 Experimental Results
Performance comparison with baselines. Main performance comparison results are shown in Tab.˜2. Firstly, zero-shot prompting yields low performance: without any task-specific adaptation Qwen-2.5-3B-Instruct achieves only 0.32% exact-action accuracy, confirming that long-horizon web behavior cannot be recovered from generic instruction tuning alone. Second, RL with sparse binary rewards on their own still fail to give the agent meaningful guidance. When we train the policy from scratch under this signal, it reaches only 1.01% exact-match action accuracy and 6.17% type accuracy. Third, a straightforward round of SFT is more effective, boosting performance to 16.76% exact match accuracy and 22.25% type accuracy. This confirms that dense, teacher-forced trajectories are crucial for injecting the structural knowledge (context $\rightarrow$ rationale $\rightarrow$ action) and illustrating the shape of the long-text fields (button labels or search queries) that the binary signal alone cannot convey. Fourth, appending an additional binary-reward RL phase after SFT delivers only mixed results: exact-match action accuracy actually slips to 16.55%, while type-level F1 rises to 28.07%. The agent thus learns to guess the coarse intent better, but it still struggles to reproduce the long-text values that drive the exact-match metric. In other words, the policy becomes better at guessing the coarse action type, but slightly worse at reproducing the fine-grained, long-text values required
@@
for an exact match. Lacking finer-grained credit assignment, the binary objective cannot push the model beyond what SFT already achieves and in some respects even pulls it backwards. Training with this objective is furthermore prone to instability and converges substantially more slowly than optimization with richer, structured rewards. Our proposed Shop-R1 framework closes most of this gap. By combining hierarchical rewards, self-certainty signals, format rewards and difficulty-aware scaling, it delivers 27.72% exact-action accuracy (+65% relative to SFT) and pushes action-type accuracy and F1 to 36.40% and 31.28%, respectively. The simultaneous rise of both the coarse (type-level) and fine-grained (exact-match) metrics indicates that Shop-R1 not only identifies the correct intent more often, but also reproduces the long-text values (button labels, text queries) with higher fidelity. More performance comparison results on OPeRA dataset (Wang et al., 2025c) can be found in Appx. E.
@@
Models Settings Exact Action Acc. Per Action Type Action Type Acc. Per Action Type click type_and_submit terminate click type_and_submit terminate Qwen-2.5-3B-Instruct Zero-shot prompting 0.58% 0.15% 0.00% 38.7% 1.62% 0.00% SFT 4.93% 3.84% 49.80% 8.55% 15.36% 49.80% SFT + RL (Binary) 8.12% 3.25% 45.51% 17.25% 13.88% 45.51% Shop-R1 (Ours) 7.39% 7.53% 81.84% 10.29% 28.66% 81.84% Qwen-2.5-1.5B-Instruct Zero-shot prompting 1.01% 0.15% 0.39% 10.00% 0.44% 0.39% SFT 4.49% 7.83% 23.44% 15.07% 32.35% 23.44% Shop-R1 (Ours) 3.62% 8.12% 72.85% 6.52% 34.12% 72.85% Qwen-2.5-0.5B-Instruct Zero-shot prompting 0.43% 0.15% 24.02% 12.90% 4.43% 24.02% SFT 3.19% 7.68% 21.88% 5.94% 26.59% 21.88% Shop-R1 (Ours) 0.72% 3.99% 97.07% 1.01% 17.87% 97.07%
As shown in Tab.˜3, we decompose accuracy by action type. Zero-shot prompting shows the classic “intent–content” split. For example, it can guess that a ‘click’ is needed (38.7% type accuracy) yet almost never names the exact UI target (0.58 % exact). Even if SFT can boost the performance but the gain remains uneven, which suggests that teacher forcing alone does not give the model enough credit assignment signal for predicting high-entropy arguments such as search queries. Appending a sparse binary RL phase after SFT still fails to boost these harder text-generation cases. Shop-R1 reshapes those incentives, higher exact match accuracy is achieved, indicating that the model is no longer satisfied with merely selecting correct type but is learning to identify the correct widget and query text as well. To be summarized, dense and structured feedback is essential: it overcomes the no-reward plateau, and makes reward hacking uneconomical.
@@ 4.3 Ablation Study And Analysis
Model size. Tab.˜2 and Tab.˜3 reveal a consistent scaling trend. In the zero-shot regime, the 3B backbone already outperforms its 1.5B and 0.5B counterparts by a factor of $\times 4\sim 5$ on coarse action–type accuracy, confirming that larger models possess stronger out-of-the-box priors for human behavior simulation at the setting of website shopping. After SFT, all sizes gain, yet the improvement is more pronounced for the two smaller backbones than for the 3B model, suggesting that demonstration learning compensates for limited capacity. Shop-R1 lifts every backbone to its best operating point, but the shape of the gains differs by scale. The 3B variant reaches the highest overall numbers while distributing its improvements evenly across the two complicated action types. By contrast, the 0.5B model achieves a comparable headline exact match accuracy (27.72%) almost entirely by over-predicting the easiest ‘terminate’ action (97.07% exact) and largely ignoring the more semantically demanding classes. The 1.5B backbone sits in between, recovering moderate fidelity on ‘click’ and ‘type_and_submit’ while retaining a strong but not overwhelming bias toward ‘terminate’. In short, scaling primarily augments the model’s ability to handle long-text, high-entropy actions; smaller networks can still match aggregate accuracy by exploiting the high-reward termination branch, but they do so at the cost of behavioral diversity. These findings underscore that, although Shop-R1 markedly mitigates capacity limitations, genuine mastery of simulation-oriented web-shopping action prediction tasks continues to benefit from larger backbones.
@@
Sampling temperature. Fig.˜2 shows that Shop-R1 is robust to sampling temperature, yet the three evaluation metrics react in distinct ways that reveal how temperature sampling propagates through the decision hierarchy. Action-type accuracy remains almost constant ($36\%$) across the entire temperature sweep because this metric aggregates all predictions: small mis-classifications in one direction are largely offset by fixes in another, leaving the overall hit rate unchanged. By contrast, the F1 score declines steadily ($31.28\%\rightarrow 28.36\%$) as temperature rises; class-averaging penalizes any asymmetric increase in confusion. Interestingly, a modest boost from the default $\tau=0.6$ to $\tau=0.7$ improves exact-match accuracy to its peak of $28.63\%$: a trace of stochasticity helps the generated response escape local maxima and occasionally assemble the full long-text argument that greedy decoding would miss. When $\tau>0.8$ the added entropy no longer uncovers new correct completions; instead it corrupts fine-grained fields faster than it fixes them, so exact-match plateaus while F1 continues to erode. This pattern is expected since the SFT stage already anchors the model to dataset-specific behavior, privileging faithful simulation over creativity. Taken together, these trends indicate that temperatures in the 0.6–0.8 band offer the best trade-off, preserving robust intent classification, maximizing strict exact-match, and avoiding the metric degradation that emerges once the sampler becomes overly exploratory.
@@
Model Training Scheme Components Exact Action Action Type SFT Format Reward Rationale Reward Reward Scale Action Reward Acc. Acc. F1 Qwen-2.5-3B-Instruct hierarchical 4.63% 36.56% 21.92% hierarchical 2.87% 3.19% 5.04% hierarchical 26.93% 37.25% 33.74% hierarchical 27.83% 27.20% 11.70% binary 27.04% 27.46% 12.11% hierarchical 27.72% 36.40% 31.28%
Training component. Tab.˜4 makes clear that every element of Shop-R1 addresses a different pathology. Removing the SFT warm-start cripples the agent: despite having all RL signals, exact-match drops to 4.63%, underscoring that a supervised prior is indispensable for learning the shape of long-text arguments. Omitting the format reward is even more destructive, where exact accuracy plunges to 2.87% and type-level metrics fall below 6% since unparseable JSON outputs earn zero credit, starving the learner of gradient signal. When the self-certainty (rationale) reward is ablated, coarse intent prediction remains strong but exact-match lags the full system by 0.8%, indicating that explicit feedback on the generated rationales mainly tightens the long-text portion of an action rather than its top-level label. Disabling the difficulty-aware reward scaling or reverting to a binary action reward leads to a different failure mode: the model still attains around 27% exact accuracy, yet type-level F1 degrades to 11–12%. Inspection shows that, without either scaling or hierarchical credit, the agent gravitates toward the easy high-reward ‘terminate’ action and rarely ventures into harder ‘click’ or ‘type_and_submit’ cases, which is a classic reward-hacking pattern. The full configuration combines all signals and delivers the best balance demonstrating that each component is necessary: SFT injects linguistic priors, the format reward safeguards parsability, the self-certainty term refines long-text precision, and hierarchical difficulty-scaled rewards prevent degenerate policies while promoting fine-grained action fidelity.
Whole-session v.s. latest-step context. Tab.˜5 isolates the impact of including the simplified HTML of each visited page in the action history. Removing this structural cue slashes exact-match accuracy from 27.72% to 14.74%, a nearly 50% relative loss, while coarse action-type accuracy drops more modestly. The sharp divergence indicates that, although the model can still infer which action type of interaction is likely next from the dialogue trace alone, it struggles to generate the fine-grained arguments, the precise button label or query string, without access to the page’s detailed context. Interestingly, the class-balanced F1 score rises slightly, suggesting that the only latest-step context
Table 5: Comparison of model performance when using either the whole-session context or only the latest-step context as input.
Context Settings Exact Action Action Type Acc. Acc. F1 whole-session 27.72% 36.40% 31.28% latest-step 14.74% 30.46% 33.48%
@@ 5 Conclusion
LLM Usage Statement
Large Language Models (LLMs) were used solely as writing assistants. Specifically, the authors drafted the initial paragraphs of the paper, and then employed an LLM to polish the language for clarity and readability. The final paragraphs were obtained after multiple rounds of human–LLM interaction, with the authors carefully reviewing, editing, and approving all content. The research ideas, experimental design, implementation, and analysis were entirely conducted by the authors.
@@
Appendix A Theoretical Analysis: Negative Rewards vs. Zero Rewards in GRPO
A common concern in reinforcement learning is that shifting rewards by a constant value can introduce biases related to episode length (e.g., “survival” or “suicide” incentives) when optimizing for discounted cumulative returns ($G_{t}=\sum\gamma^{k}r_{t+k}$). However, in the context of Group Relative Policy Optimization (GRPO), we demonstrate that the optimization landscape is invariant to affine transformations of the reward function. Consequently, setting the failure penalty to zero or a negative value yields mathematically identical gradients.
A.1 Proof of Affine Invariance
Let $\mathcal{R}=\{r_{1},r_{2},\dots,r_{G}\}$ be a set of rewards generated by a group of $G$ outputs for a single input query. GRPO computes the advantage $A_{i}$ for the $i$-th output via group-wise standardization:
$A_{i}=\frac{r_{i}-\mu}{\sigma}$ (A1)
where $\mu=\mathbb{E}[\mathcal{R}]$ is the group mean and $\sigma=\sqrt{\text{Var}(\mathcal{R})}$ is the group standard deviation.
Consider two reward schemes: Scheme A with rewards $r$ (e.g., failure penalty 0) and Scheme B with rewards $r^{\prime}$ (e.g., failure penalty $n<0$). We assume Scheme B maintains the relative order of Scheme A, meaning $r^{\prime}$ is an affine transformation of $r$:
$r^{\prime}=\alpha r+\beta,\quad\text{where }\alpha>0$ (A2)
For example, if Scheme A uses $\{1,0\}$ for success/failure and Scheme B uses $\{1,-1\}$, then $\alpha=2$ and $\beta=-1$. We now derive the advantage $A^{\prime}_{i}$ for Scheme B.
First, the new mean $\mu^{\prime}$ is:
$\mu^{\prime}=\mathbb{E}[\alpha r+\beta]=\alpha\mathbb{E}[r]+\beta=\alpha\mu+\beta$ (A3)
Second, the new standard deviation $\sigma^{\prime}$ is:
$\sigma^{\prime}=\sqrt{\text{Var}(\alpha r+\beta)}=\sqrt{\alpha^{2}\text{Var}(r)}=|\alpha|\sigma=\alpha\sigma$ (A4)
Substituting $\mu^{\prime}$ and $\sigma^{\prime}$ into the advantage formula:
$A^{\prime}_{i}=\frac{r^{\prime}_{i}-\mu^{\prime}}{\sigma^{\prime}}=\frac{(\alpha r_{i}+\beta)-(\alpha\mu+\beta)}{\alpha\sigma}=\frac{\alpha(r_{i}-\mu)}{\alpha\sigma}=\frac{r_{i}-\mu}{\sigma}=A_{i}$ (A5)
Since $A^{\prime}_{i}=A_{i}$, the advantage values—and thus the policy gradients—are identical.
A.2 Numerical Example
To illustrate this property with a realistic, unbalanced distribution, consider a group of $G=4$ samples containing 1 success and 3 failures.
  • Scheme A (Zero Penalty): Rewards are $\{1,0,0,0\}$.
    • $\mu=0.25$, $\sigma\approx 0.433$.
    • $A_{\text{success}}=\frac{1-0.25}{0.433}\approx\mathbf{1.73}$.
    • $A_{\text{failure}}=\frac{0-0.25}{0.433}\approx\mathbf{-0.57}$.
  • Scheme B (Negative Penalty): Rewards are $\{1,-1,-1,-1\}$.
    • $\mu^{\prime}=-0.5$, $\sigma^{\prime}\approx 0.866$ (Note that $\sigma^{\prime}=2\sigma$).
    • $A^{\prime}_{\text{success}}=\frac{1-(-0.5)}{0.866}\approx\mathbf{1.73}$.
    • $A^{\prime}_{\text{failure}}=\frac{-1-(-0.5)}{0.866}\approx\mathbf{-0.57}$.
As shown, the computed advantages are identical. This confirms that within the GRPO framework, the choice between zero penalty and negative penalty does not affect the optimization trajectory, provided the relative ranking of rewards is preserved.
Appendix B System Prompt and Action Space